OpenStudy (anonymous):

differentiate ln^2(t+1)

OpenStudy (anonymous):

Ln(x) will differentiate into \[\frac{ \frac{ dy }{dx} }{x }\] In this case our X=t+1 dy/dx=1 Just put these figures in for your answer. Hope that helps you out.

OpenStudy (anonymous):

the answer is 2ln(t+1)/t+1 i dont understand how to differentiate a squared log

OpenStudy (anonymous):

My bad didn't see the ^2

zepdrix (zepdrix):

The notation is a little hard to get used to. It's the same thing we do with trig functions.\[\Large\rm \ln^2(x)=\left[\ln x\right]^2\]The whole thing is being squared. So your `outermost function` is the square.

zepdrix (zepdrix):

Chain rule like you normally would, but deal with the square `before` the log.

OpenStudy (anonymous):

I see you have this, hope he helps you out HarryPotter4

OpenStudy (anonymous):

thank you guys!! i get it :)

zepdrix (zepdrix):

Got it? c: Ok cool.